OpenStudy (anonymous):

y/z ./. z/t =1 for y

OpenStudy (anonymous):

solve each formula for the given letter

OpenStudy (anonymous):

y/z divide z/t = 1 for y

OpenStudy (anonymous):

invert z/t and multiply y/z * t/z = yt / z^2 = 1 yt = z^2 y = z^2 / t